OpenStudy (anonymous):

The Pythagorean Theorem is only used for one type of triangle. Name the type of triangle that it is used for. Hint: the triangle is named after one of its angles The Pythagorean Theorem uses a formula. List the formula. True or False, any time that you have the type of triangle from part ‘a’ above…..if you know the length of 2 of the sides, you can use the formula from part ‘b’ to figure out the missing third side? What is the special name given to the longest side of the type of triangle in part ‘a’, and which variable in the formula from part ‘b’ is always used for this side?

OpenStudy (anonymous):

I know it is right triangle ( triangle in which one angle is 90 degrees) but what's the rest?

OpenStudy (anonymous):

Because really it's a special case of the cosine rule c^2 = a^2+b^2 - 2abcos(C) When C = 90 degrees, it is pythagoras' theorem this is also why it can't be used on any other triangle.

OpenStudy (anonymous):

The formula is a2 + b2 = c2. The answer to the true or false question is True because if you know the two sides, you can find the missing side by using those two sides. The longest side called the hypotenuse.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Should I put what I said into my answer as well?

OpenStudy (anonymous):

Yeah, sure.